# AllSpark: a multimodal spatiotemporal general model

### 31 Dec 2023 v1

@@ 4.3.4 Table
Yes
Stacked ResNet-LSTMcheng2022stacked
40.68 23.75 0.8
CBAM-CNN-Bi LSTMli2022prediction
18.9 11.2 0.94
For the Table modality, we evaluated AllSpark on the regression prediction task using the PRSAzhang2017cautionary dataset. The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section 4.2). The performance metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared (R2).
Table 7 presents the comparative results between AllSpark and baselines. It’s worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi LSTM proposed by D. Li et al.li2022prediction uses CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the Stacked ResNet-LSTM model proposed by X. Cheng et al.cheng2022stacked employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: with expert structures and without expert structures. In models without expert structures, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method in models with expert structures (CBAM-CNN-Bi LSTM). This reflects the excellent adaptability of AllSpark to the Table modality.
4.3.5 Code
Table 8: Code Document Generation with AllSpark
Expert Knowledge Method MRR
@@ 4.3.5 Code
No
BIRNN[68]
0.084 0.153
1D-CNN[68]
0.245 0.352
selfAtt[68]
0.365 0.451
NBoW[68]
0.429 0.461
RoBERTa[67]
0.625 0.606
@@
Yes
RoBERTa(Code)[67]
0.661 0.64
CodeBERT[67]
0.693 0.706
GraphCodeBERT[67]
0.732 0.711
For code modality, we evaluated the performance of AllSpark on the Code Document Generation task using the CodeSearchNet dataset. This task involves generating corresponding documents based on the provided function code. We conducted tests for both Ruby and JavaScript languages, using Mean Reciprocal Rank (MRR)guo2020graphcodebert as the evaluation metric.
Similar to the previous modalities, since MSI-AGI does not possess expert knowledge in the Code modality, we categorized the baselines into two groups: with and without expert knowledge. As shown in Table 8, AllSpark achieved the SOTA results in the group without expert knowledge and remained comparable to models with expert knowledge. This demonstrates the strong adaptability of AllSpark to the Code modality.
4.3.6 PointCloud
For pointcloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with Top-1 Accuracy as the metric. In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties like permutation invariance and symmetry in 3D point clouds. However, these structures designed for the unique priors of the modality are challenging to transfer across modalities. Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
Table 9: PointCloud classification with AllSpark
@@ 4.3.6 PointCloud
The design of AllSpark is based on a general sequence-to-sequence architecture. As shown in Table 9, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-net). It maintains comparability with the state-of-the-art model, PointGPT, which has both a modal expert structure and modal expert knowledge. This suggests that AllSpark holds significant potential for applications in the point cloud modality.
4.3.7 Trajectory
Table 10: Trajectory predicition with AllSpark
@@ 4.3.7 Trajectory
For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task. This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period. We report accuracy using Average Displacement Error (ADE) and Final Displacement Error (FDE)shi2023trajectory . Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (Ground Truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, ADE and FDE are used to measure their $L2$ distances, calculated as follows:
In Table 10, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, using a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to the SOTA model (STAR) with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This suggests that AllSpark demonstrates excellent adaptability to the trajectory modality.
4.3.8 SAR
Table 11: SAR classification with AllSpark
@@ 4.3.8 SAR
For SAR modality, AllSpark’s adaptability was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is Top1 Accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC (standard operating conditions) settings from AConvNetschen2016target . Table 11 presents the comparison results between AllSpark and the state-of-the-art model under these settings. Since AllSpark lacks SAR image modality expert knowledge and the spatial invariance prior to changes in perspective, its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%. This is comparable to the EMACH method. Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for our future research.
4.3.9 Infrared
AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human re-identification task involving RGB and infrared modalities. In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images. Table 12 presents the comparison results between AllSpark and the state-of-the-art re-identification methods, with the evaluation metric being Rank at 20 accuracy. The results indicate that AllSpark (76.31%) achieved similar accuracy compared to HSME (77.95%). While our method did not reach the highest accuracy, it has already surpassed some expert models in RGB-Infrared re-identification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, showcasing the potential of AllSpark in handling the infrared modality.
Table 12: Visible-Infrared Re-identification with AllSpark
@@ 4.3.9 Infrared
4.3.10 Graph
For the Graph modality, AllSpark’s performance was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include RMSE, MAE, and R2. Table 13 presents the comparison between AllSpark and the state-of-the-art results on the METR-LA dataset, with the baseline derived fromliu2023spatio . The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of RMSE, demonstrating AllSpark’s excellent adaptability to the Graph modality.
Table 13: Traffic prediction with AllSpark
@@ 4.3.10 Graph
4.3.11 Language
Table 14: Text understanding with AllSpark
@@ 4.3.11 Language
Mixed VATsachan2019revisiting ; miyato2016adversarial
@@
We tested AllSpark’s natural language processing capabilities on the IMDB datasetmaas2011learning with the task of binary sentiment classification (positive or negative). AllSpark loaded weights from the Lynx, and therefore, it can be considered to possess expert knowledge in the natural language modality. When compared with the state-of-the-art models on the IMDB dataset, as shown in Table 14, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
4.3.12 Video
For the Video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset. The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being Top1 accuracy. In Table 15, we compared AllSpark with the current state-of-the-art models. Currently, AllSpark’s adaptation to the Video modality is not optimal, showing a significant gap compared to baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark, and we trained it for only 3 epochs on the UCF101 dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective in capturing temporal information.
Table 15: Video classification with AllSpark
@@ 4.3.12 Video
4.3.13 Oblique Photography
For the Oblique Photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset. The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used is the Percentage of Accurate Grids in Total (PAG)liu2023deep , calculated by the following formula:
@@ 4.3.13 Oblique Photography
The suffixes in PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 meters, and ${PAG}_{10}$ indicates an error within 1 meter. Table 16 presents the comparison of AllSpark with popular multi-view 3D reconstruction models on the WHU-OMVS dataset. AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, it lacks the ability to design specific model structures for processing, severely restricting its performance. Adapting multi-modal large models to dense prediction tasks and optimizing the architecture is one of our future research directions.
Table 16: 3D reconstruction with AllSpark
@@
5 Discussion
5.1 Limitation
Certainly, our work has some limitations, and these will guide our future research directions:
@@ 5.1 Limitation
    Lack of interaction between different modalities. While AllSpark incorporates interaction between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks. This is primarily due to constraints such as the lack of multi-modal paired data, the inflexibility of the model architecture, and the limited richness of multi-modal interaction tasks. In the future, we plan to explore and work more extensively in these directions.
@@
    Lack of in-depth research on multi-modal alignment methods. While we introduced the basic principles of the Language as Reference Framework (LaRF) and aligned all modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method. We haven’t conducted in-depth research into the various methods of multi-modal alignment. Moreover, our testing focused on the model’s adaptability to different modalities, and we haven’t explored potential capabilities such as interactive reasoning, zero-shot learning, and multi-modal emergence that the model may possess after alignment with language. In the future, we plan to delve deeper into researching multi-modal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
@@
    Initial work. Our current work is still in its initial exploratory phase, and we haven’t carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, SAR, and video. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
@@
    Expensive Cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. Although we freeze most parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities remains one of our future research directions.
@@
    Interesting phenomenon. During our exploratory experiments, we discovered some interesting properties of large models. For instance, spatial information tends to degrade in large models, leading to collapse when performing dense prediction tasks such as segmentation and detection. Additionally, these models struggle to optimize when transferred to a small amount of downstream data. Large models often require smaller hyperparameters and are sensitive to them. These observations might partially reveal the working mechanisms of large models, and we plan to conduct more in-depth investigations into these phenomena in the future.
5.2 Potential of the LaRF
Inspired by the human cognitive system and linguistic philosophy, we propose ”Language as Reference Framework (LaRF)” as the first principle for constructing our unified multimodal model. Its foreseeable potential includes at least the following three points:
@@ 5.2 Potential of the LaRF
    Efficient Generalization of Large Models. Currently, the computational power and data scale required for training large models are rapidly expanding, and even the cost of fine-tuning large models is becoming prohibitive. Therefore, in the future, training large models for every domain is almost an impossible task. Language, however, holds the potential for achieving efficient generalization of large models, expanding them from their native domains to additional domains at minimal cost. In our proposed AllSpark, we have designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion. In theory, AllSpark can extend to any number of modalities. In the future, we will conduct more in-depth research on the impact of text prompts and lightweight parameter modules on the generalization of large models.
@@
    Interpretable Reasoning. Deep learning models have often been referred to as ”black boxes,” indicating that the reasoning process of these models is invisible and challenging to interpret. Some research on the interpretability of deep learning models is often based on complex mathematical models and numerous assumptions, greatly limiting the practical applications of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality is crucial. However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to directly use natural language for outputting reasoning chains and justifications.
@@
    Transition from end-to-end to interactive paradigm. The end-to-end paradigm refers to the learning approach where the model takes input and directly outputs results. In recent years, the end-to-end paradigm has become increasingly popular due to its simple and clear architecture and excellent performance. However, it also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues. A LaRF-based architecture has the potential to achieve an interactive paradigm, where users input raw data and corresponding text prompts, and the model automatically performs relevant operations based on the prompts. Users can even iteratively adjust the text prompts based on the results. Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to the end-to-end paradigm.
6 Conclusion

### 25 Feb 2024 v2

@@ IV-C4 Table
Yes Stacked ResNet-LSTM[74] 40.68 23.75 0.8
CBAM-CNN-Bi-LSTM[75] 18.9 11.2 0.94
Table VI presents the comparative results between AllSpark and the baselines. It is worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi-LSTM proposed by D. Li et al.[75] used a CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the stacked ResNet-LSTM model proposed by X. Cheng et al.[74] employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: those with an expert architecture and those without an expert architecture. Among the models without expert architecture, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method for models with expert architecture (CBAM-CNN-Bi-LSTM). This reflects the excellent adaptability of AllSpark to the table modality.
IV-C5 Code
For the code modality, we evaluated the performance of AllSpark on the code document generation task using the CodeSearchNet dataset. This task involves generating corresponding documents based on the provided function code. We conducted tests for both the Ruby and JavaScript languages using the mean reciprocal rank (MRR)[76] as the evaluation metric.
TABLE VII: code document generation with allspark
Expert Know. Method MRR
@@ IV-C5 Code
No BIRNN[68] 0.084 0.153
1D-CNN[68] 0.245 0.352
selfAtt[68] 0.365 0.451
NBoW[68] 0.429 0.461
RoBERTa[67] 0.625 0.606
@@
Yes RoBERTa(Code)[67] 0.661 0.64
CodeBERT[67] 0.693 0.706
GraphCodeBERT[67] 0.732 0.711
Like in the previous modalities, since MSI-AGI does not possess expert knowledge of the code modality, we categorized the baselines into two groups: those with and without expert knowledge. As shown in Table VII, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge. This finding demonstrates the strong adaptability of AllSpark to the code modality.
IV-C6 PointCloud
For the PointCloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with the top-1 accuracy as the metric. In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties such as permutation invariance and symmetry in 3D point clouds. However, these structures designed for the unique priors of the modality are challenging to transfer across modalities. Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
TABLE VIII: pointcloud classification with allspark
@@ IV-C6 PointCloud
The AllSpark module is designed based on a general sequence-to-sequence architecture. As shown in Table VIII, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net). This approach maintains comparability with the state-of-the-art PointGPT model, which has both a modal expert structure and modal expert knowledge. This finding suggested that AllSpark has significant potential for applications in the point cloud modality.
IV-C7 Trajectory
For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task. This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period. We report accuracy using the average displacement error (ADE) and final displacement error (FDE)[37]. Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (ground truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, the ADE and FDE are used to measure their $L2$ distances, calculated as follows:
In Table IX, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, which uses a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to that of the SOTA model (STAR), with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This finding suggested that AllSpark demonstrated excellent adaptability to the trajectory modality.
TABLE IX: trajectory prediction with allspark
@@ IV-C7 Trajectory
IV-C8 SAR
For the SAR modality, the adaptability of AllSpark was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is the top-1 accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC settings from AConvNets[22]. Table X presents the comparison results between AllSpark and the state-of-the-art model under these settings. Since AllSpark lacks expert knowledge of SAR image modalities and spatial invariance prior to changes in perspective; its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%. This approach is comparable to the EMACH method. Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for future research.
TABLE X: sar classification with allspark
@@ IV-C8 SAR
IV-C9 Infrared
AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human reidentification task involving RGB and infrared modalities. In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images. Table XI compares AllSpark and the other state-of-the-art reidentification methods, with the evaluation metric being rank at an accuracy of 20. The results indicate that AllSpark (76.31%) achieved similar accuracy to that of HSME (77.95%). While our method did not reach the highest accuracy, it surpassed some expert models in RGB-infrared reidentification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, demonstrating the potential of AllSpark in handling the infrared modality.
TABLE XI: visible-infrared reidentification with allspark
@@ IV-C9 Infrared
IV-C10 Graph
TABLE XII: traffic prediction with allspark
@@
For the graph modality, the performance of AllSpark was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include the RMSE, MAE, and R2. Table XII compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[39]. The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of the RMSE, demonstrating its excellent adaptability to the graph modality.
IV-C11 Language
We tested AllSpark’s natural language processing capabilities on the IMDB dataset[43] with the task of binary sentiment classification (positive or negative). AllSpark loads weights from the Lynx; therefore, it can be considered to possess expert knowledge in the natural language modality. Compared with the SOTA models on the IMDB dataset, as shown in Table XIII, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
TABLE XIII: text understanding with allspark
@@
Mixed VAT[94, 95]
@@
IV-C12 Video
For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset. The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy. In Table XIV, we compare AllSpark with the current state-of-the-art models. Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark. We trained it for only 3 epochs on the UCF101 dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
TABLE XIV: video classification with allspark
@@
IV-C13 Oblique Photography
For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset. The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used was the percentage of accurate grids in total (PAG)[42], calculated by the following formula:
@@
The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre. Table XV compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset. AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance. Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
TABLE XV: 3D reconstruction with allspark
@@
V Discussion
V-A Limitations
Certainly, our work has several limitations, which will guide our future research directions:
@@ V-A Limitations
    Lack of interaction between different modalities. While AllSpark incorporates interactions between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks. This is primarily due to constraints such as the lack of multimodal paired data, the inflexibility of the model architecture, and the limited richness of multimodal interaction tasks. In the future, we plan to explore and work more extensively in these directions.
@@
    There is a lack of in-depth research on multimodal alignment methods. While we introduced the basic principles of the LaRF and aligned all the modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method. We have not conducted in-depth research into the various methods of multimodal alignment. Moreover, our testing focused on the model’s adaptability to different modalities, and we have not explored potential capabilities such as interactive reasoning, zero-shot learning, or multimodal emergence that the model may possess after alignment with language. In the future, we plan to delve deeper into researching multimodal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
@@
    Initial work. Our current work is still in its initial exploratory phase, and we have not carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain imaging modalities, such as oblique photography, SAR, and video. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
@@
    Expensive cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. Although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities is one of our future research directions.
@@
    Interesting phenomenon. During our exploratory experiments, we discovered several interesting properties of large models. For instance, spatial information tends to degrade in large models, leading to collapse when performing dense prediction tasks such as segmentation and detection. Additionally, these models struggle to optimize when transferred to a small quantity of downstream data. Large models often require smaller hyperparameters and are sensitive to them. These observations might partially reveal the working mechanisms of large models, and we plan to conduct additional in-depth investigations into these phenomena in the future.
V-B Potential of the LaRF
Inspired by the human cognitive system and linguistic philosophy, we propose the ”LaRF” as the first principle for constructing our unified multimodal model. Its foreseeable potential includes at least the following three points:
@@ V-B Potential of the LaRF
    Efficient Generalization of Large Models. Currently, the computational power and data scale required for training large models are rapidly expanding, and even the cost of fine-tuning large models is becoming prohibitive. Therefore, in the future, training large models for every domain will be almost impossible. Language, however, holds the potential for achieving efficient generalization of large models, expanding them from their native domains to additional domains at minimal cost. With our proposed AllSpark, we designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion. In theory, AllSpark, built on the LaRF principle, can be extended to arbitrary modalities. In the future, we will conduct more in-depth research on the impact of text prompts and lightweight parameter modules on the generalization of large models.
@@
    Interpretable Reasoning. Deep learning models have often been referred to as ”black boxes,” indicating that the reasoning process of these models is invisible and challenging to interpret. Research on the interpretability of deep learning models often relies on complex mathematical models and numerous assumptions, greatly limiting the practical application of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality are crucial. However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to use natural language directly for outputting reasoning chains and justifications.
@@
    Transition from an end-to-end to an interactive paradigm. The end-to-end paradigm refers to the learning approach where the model takes input and directly outputs results. In recent years, the end-to-end paradigm has become increasingly popular due to its simple and clear architecture and excellent performance. However, this approach also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues. A LaRF-based architecture has the potential to achieve an interactive paradigm in which users input raw data and corresponding text prompts and the model automatically performs relevant operations based on the prompts. Users can even iteratively adjust the text prompts based on the results. Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to those of the end-to-end paradigm.
VI Conclusion